Interactive LearningWare 27.2 reviews the concepts that are important in this problem. A uniform layer of water (n = 1.33) lies on a glass plate (n = 1.52). Light shines perpendicularly on the layer. Because of constructive interference, the layer looks maximally bright when the wavelength of the light is 434 nm in vacuum and also when it is 651 in vacuum. (a) Obtain the minimum thickness of the film. (b) Assuming that the film has the minimum thickness and that the visible spectrum extends from 380 to 750 nm, determine the visible wavelength (in vacuum) for which the film appears completely dark.
Answer #1

6 The Condino fa Constructive Inkeference at - m?file = m( t = 2 ( 2 ) for uzunm = m (434) 6 forbsium t = ml (651) - = ③ (434